# Sparse Deformable Mamba for Hyperspectral Image Classification

### v1

@@ Abstract
Although the recent Mamba models significantly improve hyperspectral image (HSI) classification, one critical challenge is caused by the difficulty to build the Mamba sequence efficiently and effectively. This paper presents a Sparse Deformable Mamba (SDMamba) approach for enhanced HSI classification, with the following contributions. First, to enhance Mamba sequence, an efficient Sparse Deformable Sequencing (SDS) approach is designed to adaptively learn the ”optimal” sequence, leading to sparse and deformable Mamba sequence with increased detail preservation and decreased computations. Second, to boost spatial-spectral feature learning, based on SDS, a Sparse Deformable Spatial Mamba Module (SDSpaM) and a Sparse Deformable Spectral Mamba Module (SDSpeM) are designed for tailored modeling of the spatial information spectral information. Last, to improve the fusion of SDSpaM and SDSpeM, an attention based feature fusion approach is designed to integrate the outputs of the SDSpaM and SDSpeM. The proposed method is tested on several benchmark datasets with many state-of-the-art approaches, demonstrating that the proposed approach can achieve higher accuracy, faster speed, and better detail small-class preservation capability.
@@ I Introduction
In previous decades, various approaches have been proposed for feature learning from HSI. For example, principal component analysis (PCA) [1] and independent component analysis (ICA) [2] have been used to extract compact spectral features from HSI. A morphological approach [3] has also been used to enhance the spatial feature extraction. However, these approaches are feature engineering approaches that cannot fully capture the discriminative HSI information in an adaptive manner. Deep learning approaches have been widely used for improving HSI feature extraction. Convolutional neural networks (CNN) [4] improve the learning of HSI spatial-spectral features, but they have limitations in terms of strong inductive bias and locality to capture the long-range spatial-spectral correlation effect in HSI. Transformers are more flexible and adaptive to longer-range spatial context [5], but they require a large attention matrix and thereby big computational cost. Comparing with Transformers, the Mamba models, due to their token sequencing, can greatly reduce computations while maintaining the long-range modeling capacity. Hence, many Mamba approaches [6] have been proposed for improving HSI classification.
However, for Mamba models [7], one critical issue is how to build the sequence in a compact and efficient manner. First, compactness is essential for enabling longer-sequence learning and reducing computational cost and redundancy in tokens. Second, the order of tokens in Mamba sequence is critical for improving Mamba’s modeling capacity and overcoming the correlation vanishing issue. Therefore, instead of scanning tokens in a predefined deterministic manner, defining token sequence in an adaptive and learnable manner is critical for boosting Mamba modeling capacity.
Figure 1: Illustration of our proposed sparse deformable sequencing (SDS) for improving Mamba. Various classical scanning approaches in (a)-(h) are dense and predefined, because they use all tokens in a deterministic manner, causing potential redundancy and rigidity to adapt to data characteristics, unnecessary computation cost, and difficulty in selecting scanning approaches. In contrast, our SDS approach in (i) can identify and sequence limited number of relevant tokens in a learnable and adaptive manner, leading to sparse and deformable sequence patterns that can reduce redundancy, rigidity and computational cost in Mamba.
@@
  • To improve compactness and efficiency in Mamba token sequencing, instead of using deterministic dense scanning approaches, a sparse deformable sequencing (SDS) approach is designed, which can identify and sequence a limited number of relevant tokens in a learnable and adaptive manner, leading to sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in Mamba. As illustrated in Figure 1, compared to various predefined scanning approaches, the proposed SDS approach tends to overcome the potential redundancy and rigidity to adapt to data characteristics, unnecessary computational cost, and difficulty selecting among various scanning approaches.
  • To improve spatial-spectral feature learning, based on SDS, a Sparse Deformable Spatial Mamba Module (SDSpaM) and a Sparse Deformable Spectral Mamba Module (SDSpeM) are designed for tailored modeling of the spatial information spectral information.
  • To enhance the fusion of SDSpaM and SDSpeM, an attention based feature fusion approach is designed to integrate the outputs of the SDSpaM and SDSpeM. The proposed method is tested several benchmark datasets with many state-of-the-art approaches, demonstrating that the proposed approach can achieve higher accuracy, faster speed, and better detail small-class preservation capability.
@@ II-A Overview of Sparse Deformable Mamba (SDMamba) Model
Figure 2 displays the architecture of the proposed SDMamba model. As we can see, the proposed SDMamba model inputs $\mathit{X_{j}}$, a $H\times W\times 200$ data cube, with $W$, $H$ and $200$ being respectively the width, height and number of channels. Taking Indian Pines dataset as an example, $H=9$, $W=9$, and there are 200 channels. After the SDSpaM and the SDSpeM modules and the attention fusion module, it outputs the label of the center pixel of this data cube, i.e., $\mathit{L_{j}}$, a $K\times 1$ tensor, where $K$ is the number of classes. The key building blocks are sparse deformable sequencing (SDS), SDSpaM, SDSpeM and attention fusion module, which are illustrated in the subsequent subsections.
Figure 2: The proposed SDMamab is sparse, because the input sequences to the Mamba Blocks, i.e., $\overline{\boldsymbol{\mathit{Z}}}_{j}(5\ tokens)$ and $\overline{\boldsymbol{\mathit{A}}}_{j}(3\ tokens)$ have much less number of tokens than $\boldsymbol{\mathit{Z}}_{j}(HW\ tokens)$ and $\boldsymbol{\mathit{A}}_{j}(C\ tokens)$ respectively. In SDMamba, the order of tokens in $\overline{\boldsymbol{\mathit{Z}}}_{j}$ and $\overline{\boldsymbol{\mathit{A}}}_{j}$ are deformable and learnable, because two adaptive attention matrices, i.e., $SparseSpatialAttn$ and $SparseSpectralAttn$, are used to sort the tokens according to their relevance and identify and use a few most relevant tokens. So, the order of tokens will change if $SparseSpatialAttn$ and $SparseSpectralAttn$ change. Therefore, the proposed SDMamba approach has sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in classical Mamba models.
@@ II-B Sparse Deformable Sequencing (SDS)
  • Instead of using all tokens as conducted in classical scanning approaches in Figure 1, how to reduce the potential redundancy and computational cost by building sparse and compact sequence using the most relevant tokens?
  • Instead of defining the sequence in a deterministic predefined approach, how to adaptively build a deformable sequence in a learnable manner to overcome the rigidity of classical scanning approaches?
@@
Second, in SDMamba, the order of tokens in $\overline{\boldsymbol{\mathit{Z}}}_{j}$ and $\overline{\boldsymbol{\mathit{A}}}_{j}$ are deformable and learnable, because two adaptive attention matrices, i.e., $SparseSpatialAttn$ and $SparseSpectralAttn$, are used to sort the tokens according to their relevance and identify and use a few most relevant tokens. So, the order of tokens will change if $SparseSpatialAttn$ and $SparseSpectralAttn$ change. Therefore, the proposed SDMamba approach has sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in classical Mamba models.
@@ II-C Sparse Deformable Spatial Mamba Module (SDSpaM)
In Figure 2, the SDSpaM module takes as input $\mathit{Z_{j}}$, which has a total of $H*W$ tokens, with each token being a $C\times 1$ vector. Instead of using $\mathit{Z_{j}}\ (H*W\ tokens)$ directly as input to the $MambaBlock$, we generate a Sparse Deformable Spatial Sequence, denoted by $\overline{\boldsymbol{\mathit{Z}}}_{j}\ (5\ tokens)$ to feed the $MambaBlock$, to reduce the potential redundancy and computational cost. $SparseSpatialAttn$, a $H\times W$ matrix, is used to identify these 5 tokens, and sort these tokens according to their relevance. So, the sequence is deformable and learnable, because $SparseSpatialAttn$ is adaptively learned from the data. The output of $MambaBlock$, denoted by $\underline{\boldsymbol{\mathit{Z}}}_{j}$, is scattered into the spatial dimensions of $\mathit{Z_{j}}$, which serves as a residual skip connection that is commonly used in neural network architectures.
A stem layer is used to get the initial feature:
@@ II-D Sparse Deformable Spectral Mamba Module (SDSpeM)
In Figure 2, the SDSpeM module also takes as input $\mathit{A_{j}}$, which has a total of $C$ tokens, with each token being a $H*W\times 1$ vector. Instead of using $\mathit{A_{j}}\ (C\ tokens)$ directly as input to the $MambaBlock$, we generate a Sparse Deformable Spectral Sequence, denoted by $\overline{\boldsymbol{\mathit{A}}}_{j}\ (3\ tokens)$ to feed the $MambaBlock$, to reduce the potential redundancy and computational cost. $SparseSpectralAttn$, a $C\times 1$ vector, is used to identify these 3 tokens, and sort these tokens according to their relevance. So, the sequence is deformable and learnable, because $SparseSpectralAttn$ is adaptively learned from the data. The output of $MambaBlock$, denoted by $\underline{\boldsymbol{\mathit{A}}}_{j}$, is scattered into the spatial dimensions of $\mathit{A_{j}}$, which serves as a residual skip connection that is commonly used in neural network architectures.
We reshape $\mathbf{\mathit{Y_{j}}}$ into $\mathbf{\mathit{A_{j}}}$, a $C\times HW$ matrix, with $C$ being the number of tokens and each token being a $HW\times 1$ vector.
@@ II-E Attention Data Fusion Module
Based on the outputs ($H\times W\times 256$) of the SDSpaM module and the SDSpeM module, we design an Attention Fusion approach to leverage the attention mechanism to improve the fusion of spatial information and spectral information. First, we use the output of SDSpaM to calculate $Q$ of size $(H*W)\times 256$, and use the SDSpeM output to calculate $K$ and $K$ with the same size. An attention matrix is achieved by multiplying $Q$ with $K$, which is then used to update $V$ for achieving the fused feature map.
@@ III-B Results
Futrhermore, 2-D graphical visulization map depicting the features extracted by proposed SDMamba. Using the t-SNE, we can clearly see that test data of similar categories gather together, and intraclass variance is minimized on our proposed method.
Table III provides a comparison of different levels of sparsity ratio and floating-point operations (FLOPs) for model with or without Sparse Deformable Sequencing (SDS) approach. As we can see from the Table III, three accuracy indexes have slight changes when using different sparsity ratios, and SDS approach performs better than without sparsity operation. Additionally, SDS approach can reduce the FLOPs, which huge decrease the computational cost.
TABLE I: Quantitative performance of different classification methods in terms of OA, AA, $k$, as well as the accuracies for each class on the Indian Pines dataset with 10 % training samples. The best results are in bold.
@@
TABLE II: Quantitative performance of different classification methods in terms of OA, AA, $k$, as well as the accuracies for each class on the Pavia University dataset with 3 % training samples. The best results are in bold.
@@ IV Conclusion
In this paper, we have presented a Sparse Deformable Mamba (SDMamba) approach to enhance the HSI classification process. We have the following contributions. First, an efficient Sparse Deformable Sequencing (SDS) approach has been designed to learn the ”optimal” sequencing, which not only optimize the learning capacity of the Mamba model but also increases its efficiency with less computations. Second, the SDS approach was integrated with the spatial module and the spectral module, leading to two dedicated HSI feature learning modules, the Sparse Deformable Spatial Mamba Module (SDSpaM) and Sparse Deformable Spectral Mamba Module (SDSpeM), which are dedicated to learning the spatial context information and SDspeM focusing on the spectral information respectively. Last, a novel feature fusion approach is designed based on the attention mechanism which can efficiently integrate the output of the SDSpaM and SDSpeM for HSI classification. The proposed approach was tested on the Indian Pines and Pavia University HSI datasets in comparison with various other state-of-the-art approaches, demonstrating that our approach outperformed others in terms of both accuracy and computational cost.

### v2

@@ Abstract
Although Mamba models significantly improve hyperspectral image (HSI) classification, one critical challenge is the difficulty in building the sequence of Mamba tokens efficiently. This paper presents a Sparse Deformable Mamba (SDMamba) approach for enhanced HSI classification, with the following contributions. First, to enhance Mamba sequence, an efficient Sparse Deformable Sequencing (SDS) approach is designed to adaptively learn the ”optimal” sequence, leading to sparse and deformable Mamba sequence with increased detail preservation and decreased computations. Second, to boost spatial-spectral feature learning, based on SDS, a Sparse Deformable Spatial Mamba Module (SDSpaM) and a Sparse Deformable Spectral Mamba Module (SDSpeM) are designed for tailored modeling of the spatial information spectral information. Last, to improve the fusion of SDSpaM and SDSpeM, an attention based feature fusion approach is designed to integrate the outputs of the SDSpaM and SDSpeM. The proposed method is tested on several benchmark datasets with many state-of-the-art approaches, demonstrating that the proposed approach can achieve higher accuracy with less computation, and better detail small-class preservation capability.
@@ I Introduction
Figure 1: Illustration of our proposed sparse deformable sequencing (SDS) for improving Mamba. Various classical scanning approaches in (a)-(h) are dense and predefined, because they use all tokens in a deterministic manner, causing potential redundancy and rigidity, unnecessary computation cost, and difficulty in selecting scanning approaches. In contrast, our SDS approach in (i) can identify and sequence limited number of relevant tokens in a learnable and adaptive manner, leading to sparse and deformable sequence patterns that can reduce redundancy, rigidity and computational cost in Mamba.
@@
In previous decades, various approaches have been proposed for feature learning from HSI. For example, principal component analysis (PCA) [1] and independent component analysis (ICA) [2] have been used to extract compact spectral features from HSI. A morphological approach [3] has also been used to enhance the spatial feature extraction. However, these approaches are feature engineering approaches that cannot fully capture the discriminative HSI information in an adaptive manner. Deep learning approaches have been widely used for improving HSI feature extraction. Convolutional neural networks (CNN) [4, 5, 6] improve the learning of HSI spatial-spectral features, but they have limitations in terms of strong inductive bias and locality to capture the long-range spatial-spectral correlation effect in HSI. Transformers are more flexible and adaptive to longer-range spatial context [7, 8, 9], but they require a large attention matrix and thereby big computational cost. Comparing with Transformers, the Mamba models, due to their token sequencing, can greatly reduce computations while maintaining the long-range modeling capacity. Hence, many Mamba approaches [10, 11, 12] have been proposed for improving HSI classification.
However, for Mamba models [13], one critical issue is how to build the sequence in a compact and efficient manner. First, compactness is essential for enabling longer-sequence learning and reducing computational cost and redundancy in tokens. Second, the order of tokens in Mamba sequence is critical for improving Mamba’s modeling capacity and overcoming the correlation vanishing issue. Therefore, instead of scanning tokens in a dense, predefined, and deterministic manner, defining token sequence in a sparse, adaptive and learnable manner is critical for boosting Mamba modeling capacity.
@@
  • To improve compactness and efficiency in Mamba token sequencing, instead of using deterministic dense scanning approaches, a sparse deformable sequencing (SDS) approach is designed, which can identify and sequence a limited number of relevant tokens in a learnable and adaptive manner, leading to sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in Mamba. As illustrated in Figure 1, compared to various predefined scanning approaches, the proposed SDS approach tends to overcome the potential redundancy and rigidity, unnecessary computational cost, and the difficulty selecting among various scanning approaches.
  • To improve spatial-spectral feature learning, based on SDS, a Sparse Deformable Spatial Mamba Module (SDSpaM) and a Sparse Deformable Spectral Mamba Module (SDSpeM) are designed for tailored modeling of the spatial information spectral information respectively.
  • To enhance the fusion of SDSpaM and SDSpeM, an attention based feature fusion approach is designed to integrate the outputs of the SDSpaM and SDSpeM. The proposed method is tested several benchmark datasets with many state-of-the-art approaches, demonstrating that the proposed approach can achieve higher accuracy and better detail small-class preservation capability.
@@ II-A Overview of Sparse Deformable Mamba (SDMamba) Model
Figure 2 displays the architecture of the proposed SDMamba model. As we can see, the proposed SDMamba model ingests $\mathit{X_{j}}$, a $H\times W\times 200$ data cube, with $W$, $H$ and $200$ being respectively the width, height and number of channels. Taking Indian Pines dataset as an example, $H=9$, $W=9$, and there are 200 channels. After the SDSpaM and the SDSpeM modules and the attention fusion module, it outputs the label of the center pixel of this data cube, i.e., $\mathit{L_{j}}$, a $K\times 1$ tensor, where $K$ is the number of classes. The key building blocks are sparse deformable sequencing (SDS), SDSpaM, SDSpeM and attention fusion module, which are illustrated in the subsequent subsections.
Figure 2: The proposed SDMamab is sparse, because the input sequences to the MambaBlock, i.e., $\overline{\boldsymbol{\mathit{Z}}}_{j}(5\ tokens)$ and $\overline{\boldsymbol{\mathit{A}}}_{j}(3\ tokens)$ have much less number of tokens than $\boldsymbol{\mathit{Z}}_{j}(HW\ tokens)$ and $\boldsymbol{\mathit{A}}_{j}(C\ tokens)$ respectively. Moreover, in SDMamba, the order of tokens in $\overline{\boldsymbol{\mathit{Z}}}_{j}$ and $\overline{\boldsymbol{\mathit{A}}}_{j}$ are deformable and learnable, because two adaptive attention matrices, i.e., $SparseSpatialAttn$ and $SparseSpectralAttn$, are used to sort the tokens and identify a limited number of relevant tokens. Therefore, the proposed SDMamba approach has sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in classical Mamba models.
@@ II-B Sparse Deformable Sequencing (SDS)
  • Sparsity: Instead of using all tokens as conducted in classical scanning approaches as in Figure 1 (a)-(h), how to reduce the potential redundancy and computational cost by building sparse and compact sequence only using the most relevant tokens?
  • Learnability and Adaptability: Instead of defining the sequence in deterministic predefined manners as in Figure 1 (a)-(h), how to adaptively build a deformable sequence in a learnable manner to overcome the rigidity of classical scanning approaches?
@@
Second, in SDMamba, the order of tokens in $\overline{\boldsymbol{\mathit{Z}}}_{j}$ and $\overline{\boldsymbol{\mathit{A}}}_{j}$ are deformable and learnable, because two adaptive attention matrices, i.e., $SparseSpatialAttn$ and $SparseSpectralAttn$, are used to sort the tokens and identify a limited number of tokens.
Therefore, the proposed SDMamba approach has sparse and deformable sequence patterns that can reduce redundancy, rigidity, and computational cost in classical Mamba models.
@@ II-C Sparse Deformable Spatial Mamba Module (SDSpaM)
In Figure 2, the SDSpaM module takes as input $\mathit{Z_{j}}$, which has a total of $H*W$ tokens, with each token being a $C\times 1$ vector. Instead of using $\mathit{Z_{j}}\ (H*W\ tokens)$ directly as input to the $MambaBlock$, we generate a Sparse Deformable Spatial Sequence, denoted by $\overline{\boldsymbol{\mathit{Z}}}_{j}\ (5\ tokens)$ to feed the $MambaBlock$, to reduce the potential redundancy and computational cost. $SparseSpatialAttn$, a $H\times W$ matrix, is used to identify these 5 tokens by sorting these tokens according to their relevance. So, the sequence is deformable and learnable, because $SparseSpatialAttn$ is adaptively learned from the data. The output of $MambaBlock$, denoted by $\underline{\boldsymbol{\mathit{Z}}}_{j}$, is scattered into the spatial dimensions of $\mathit{Z_{j}}$, which serves as a residual skip connection that is commonly used in neural network architectures.
More specifically, a stem layer is used to get the initial feature:
@@ II-D Sparse Deformable Spectral Mamba Module (SDSpeM)
In Figure 2, the SDSpeM module also takes as input $\mathit{A_{j}}$, which has a total of $C$ tokens, with each token being a $H*W\times 1$ vector. Instead of using $\mathit{A_{j}}\ (C\ tokens)$ directly as input to the $MambaBlock$, we generate a Sparse Deformable Spectral Sequence, denoted by $\overline{\boldsymbol{\mathit{A}}}_{j}\ (3\ tokens)$ to feed the $MambaBlock$, to reduce the potential redundancy and computational cost.
$SparseSpectralAttn$, a $C\times 1$ vector, is used to sort these tokens and identify these 3 tokens. So, the sequence is deformable and learnable, because $SparseSpectralAttn$ is adaptively learned from the data.
The output of $MambaBlock$, denoted by $\underline{\boldsymbol{\mathit{A}}}_{j}$, is scattered into the spatial dimensions of $\mathit{A_{j}}$, which serves as a residual skip connection that is commonly used in neural network architectures.
More specifically, we reshape $\mathbf{\mathit{Y_{j}}}$ into $\mathbf{\mathit{A_{j}}}$, a $C\times HW$ matrix, with $C$ being the number of tokens and each token being a $HW\times 1$ vector.
@@ II-E Attention Data Fusion Module
Based on the outputs ($H\times W\times 256$) of the SDSpaM module and the SDSpeM module, we design an attention fusion approach to leverage the attention mechanism for improving the fusion of spatial information and spectral information. As illustrated in Figure 3, first, we use the output of SDSpaM to calculate $Q$ of size $(H*W)\times 256$, and use the SDSpeM output to calculate $K$ and $K$ with the same size. An attention matrix is achieved by multiplying $Q$ with $K$, which is then used to update $V$ for achieving the fused feature map.
@@ III-B Results
Table III shows the influence of sparsity ratios and the floating-point operations (FLOPs). As we can see, our sparse approach not only has much less FLOPS with reduced computational cost, but also higher accuracy than the classical dense token approach (i.e., w/o Sparsity in the table). In fact, using 5% of the tokens (i.e., Sparsity ratio 0.05) outperforms the classical Mamba approach on the Indian pines dataset, and achieves comparable performance on the Pavia dataset, demonstrating the significant benefits of reducing the redundancy in the Mamba sequence.
@@
Futrhermore, Figure 4 (m)-(n), and Figure 5 (m)-(n) show the TSNE visualization of features extracted by proposed SDMamba model. We can clearly see that our approach can disentangle the different classes that are hiding in the original space.
TABLE I: Quantitative performance of different classification methods in terms of OA, AA, $k$, as well as the accuracies for each class on the Indian Pines dataset with 10 % training samples. The best results are in bold and colored shadow.
@@
TABLE II: Quantitative performance of different classification methods in terms of OA, AA, $k$, as well as the accuracies for each class on the Pavia University dataset with 3 % training samples. The best results are in bold and colored shadow.
@@ IV Conclusion
In this paper, we have presented a Sparse Deformable Mamba (SDMamba) approach to enhance the HSI classification. We have the following contributions. First, an efficient Sparse Deformable Sequencing (SDS) approach has been designed to learn the ”optimal” sequencing, which not only optimize the learning capacity of the Mamba model but also increases its efficiency with less computations. Second, the SDS approach was integrated with the spatial module and the spectral module, leading to two dedicated HSI feature learning modules, i.e., the Sparse Deformable Spatial Mamba Module (SDSpaM) and Sparse Deformable Spectral Mamba Module (SDSpeM), which are dedicated to learning the spatial context information and SDspeM focusing on the spectral information respectively. Last, a novel feature fusion approach was designed based on the attention mechanism which can efficiently integrate the output of the SDSpaM and SDSpeM for HSI classification. The proposed approach was tested on the Indian Pines and Pavia University HSI datasets in comparison with various other state-of-the-art approaches, demonstrating that our approach outperformed others in terms of both accuracy and computational cost.